OpenStudy (anonymous):

what is the largest value of curvature of the f(x)=(x)^3/3

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Curvature#Curvature_of_a_graph curvature is defined as: \[k(x) = \frac{f''(x)}{(1+f'(x)^{2})^{3/2}}\] to find largest value or maximize curvature, set derivative equal to zero \[k'(x) = 0\] first find f'(x) and f''(x) then find k'(x)

OpenStudy (dumbcow):

does that help? or are you having trouble taking the derivatives

OpenStudy (anonymous):

@marsss that is completely irrelevant

OpenStudy (anonymous):

@dumbcow *curvature for z = f(x) functions is defined as that :).

OpenStudy (dumbcow):

@ee2oo , im going to respond here the denominator doesn't matter since it equals 0 .... just multiply both sides by denominator

OpenStudy (dumbcow):

@RONNCC , haha thx for correction

OpenStudy (dumbcow):

also to confirm, i think you got k'(x) right \[\rightarrow k'(x) = \frac{2(1+x^{4})^{3/2} - 12x^{4} \sqrt{1+x^{4}}}{(1+x^{4})^{3}}\]

OpenStudy (dumbcow):

\[2(1+x^{4})^{3/2} - 12x^{4} \sqrt{1+x^{4}} = 0\] \[\sqrt{1+x^{4}} [2(1+x^{4}) -12x^{4}] = 0\] \[2 -10x^{4} = 0\] \[x^{4} = \frac{1}{5}\]

OpenStudy (anonymous):

@RONNCC you are an idiot :)